Journalist Report – February 8th

Crew 274 Journalist Report 2-8-2023

Sol: 10

Author: Tony DiBernardo, Crew Journalist

Title: KURT II

Crew 274 woke up this morning at 7:30am to the song “The Final Countdown” from Europe. The crew got their surveys done, then prepared for another double EVA day utilizing KURT, the Kinetic Utilization Research Tool.

EVA Activity

Two EVAs were completed which closes out Commander Guthrie’s research on methods for rescuing incapacitated astronauts on the surface of the Moon or Mars. The study observes the performance of a medical sled in a Martian-like terrain and the use of an assistive device, rescue vest, with specialized handles for picking up the downed astronauts. The vest represents hoist points on the hard upper torso (HUT) which is the hardened structure of the spacesuit.

The intent is to learn if assistive rescue handles attached to the HUT can improve the life-saving capability of surface activities when astronauts are on interplanetary missions. NASA’s Artemis missions outline the necessity for proper suit interface and hoist connections for rescues. This study aims to use combat experiences for extracting members from the field and applying those concepts to Lunar and Martian missions.

Hab Design

Hab Specialist O’Hara continued his evaluation of habitats for this post-doctoral study and for Blue Origin by beginning his evaluation of the crew living quarters and galley.


Crew astronomers Loy and Pena observed detailed heliophysics phenomena today as they collected bulk data sets on solar dark spots and solar chromosphere prominences. Large prominences were observed wrapping around solar magnetic fields, just before solar flares were registered by NOAA blasting away from the sun. In total, 16 large solar prominences and 12 solar dark spots were observed today. These observations will be very useful for the astronomer’s analysis of solar cycle 25, how space weather implicates orbital assets, space operations planning to safeguard those assets, and the importance of having heliophysics astronomers on-site at Martian Habitats. On top of these observations, 80,000 images were captured, and 20,000 solar spots and prominences were stacked, measured, and analyzed.


65 separate 45-second long exposure images were collected from the MLCRCOS-16 New Mexico Observatory. These images are being separately captured through visual, infrared, and blue filters, assisting Astronomers Pena and Loy’s science on HADS variable star research. 15 images were stacked, analyzed, and compared to previous images of this star system, HADS Variable Star V0799 AUR, to provide an assessment of this variable star’s brightness variability period and cycle, as well as provide a true-brightness standard to accurately measure distances in the immediate vicinity of HADS V0799 AUR’s stellar neighborhood.

EVA Report – February 8th



Crew 274 Sol 9 EVA Report 02-08-2023

EVA #: 10

Author: Sarah “Ceres” Guthrie, Commander

Purpose of EVA: The purpose of this EVAs is to test the adaptive rescue vest on KURT (test dummy) with the rescue sled hilled and rocky surfaces. This location was chosen for the scenario where incapacitated astronaut would have fallen into a ravine that can not be accessed by a rover. The crew will save “KURT” from the river bed at the base of the hill and return him to the rover parked on Cow Dung Road (N4251500, E518800). NOTE: More time is requested (30min+) as the trek through the ravine took more time than anticipated in EVA #9.

Start time: 10:00am

End time: 12:00pm

Narrative: Additional crew members performed an EVA with the adaptive rescue vests with a medical sled, at the ravine at Gateway of Candor. They dragged KURT (~80 lbs) to the destination via a medical sled. The sled continues to perform well over the various terrain features (soft riverbed and rock), however it did develop a small hole caused by abrasion. The crew also simulated an incapacitated astronaut injury using KURT (a dummy) who fell from the hill, by placing him on the side of the hill slope. They tested the rescue vest, simulating a pressurized suit stance, lifting KURT at different points. Each member of the crew performed a single-person and two-member rescue configuration bringing him to the top of the hill and down different slopes via the lifting KURT at different connection points and placing KURT on the sled. They then practiced single-person and two-member carry of the sled and walked back to the rover. This test was to evaluate the performance of the sled over Martian like terrain, which was not done in previous analogs as the terrain provides insight to the challenges of performing a rescue. The rescue vest improved the ease of lifting KURT from the surface which was engineered from lessons learned in past analog settings.

Destination: Gateway of Candor

Coordinates (use UTM WSG 84): N4251500, E519500

EVA Participants: Sarah Guthrie (Crew Commander), Tyler Hines (Greenhab), Alexis Lojek (Crew Engineer), Tony DiBernardo (Journalist)

Road(s) and routes per MDRS Map: EVA #10 traveled north on Cow Dung Road to Gateway to Candor.

Mode of travel: Walking and Rovers (Spirit & Opportunity)

Astronomy Report – February 8th



Astronomy Report

Name: Salina Peña and Noah Loy

Crew: 274

Date: 02/08/2023


Robotic Telescope Requested (Choose one) MDRS-14 MDRS-WF

Objects Viewed: Star/230208/variable star

Problems Encountered: None. Used the MLC RCOS 16 for observations.


Solar Features Observed: Sun/230208/solar prominence


Problems Encountered: None

GreenHab Report – February 8th



GreenHab Officer: Tyler Hines

Environmental control: heater

Average temperatures: 75.7 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: 1.056 gallons

Water in Blue Tank: 59.89 gallons

Time(s) of watering for crops: 1:01 pm

Changes to crops: N/A

Narrative: Continued to nominally maintain and water crops and plants as part of daily GreenHab operations. In addition to a final harvest list being prepared ahead of the final days of crew operations and meal planning, another effort to trim plants and crops along with general cleaning of the GreenHab was also performed.

Harvest: N/A

Support/supplies needed: N/A

Operations Report – February 8th



Crew 274 Operations Report 2/8/2023

SOL: 10

Name of person filing report: Alexis Lojek

Non-nominal systems: Toilet. Robotic Observatory.

Notes on non-nominal systems: Mission Support will install a new toilet holding tank between crews. Robotic observatory is closed until further notice.


Spirit rover used: Yes

Hours: 208.5

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.5

Beginning charge: 100%

Ending charge: 76%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 220.7

Beginning charge: 100%

Ending charge: 80%

Currently charging: No

Perseverance rover used: Yes

Hours: 255.4

Beginning charge: 100%

Ending charge: 71%

Currently charging: No

General notes on rovers: Rovers Spirit and Opportunity were used for the morning EVA and plugged in when they returned, as it was a cloudless day and before noon. Curiosity and Perseverance were utilized for the afternoon EVA, but returned after noon, so were not plugged.

Summary of Hab operations: Toilet tank was emptied this morning and habitat is in a nominal condition.

WATER USE: 36 gallons as of 5:30pm

Water (static tank): 278 gallons remaining

Static tank pipe heater (On or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits are in nominal condition and charging. Radios were charged yesterday and are off chargers. All conditions normal.

Summary of GreenHab operations: GHO Hines watered plants from 1:01 – 1:32 pm.

WATER USE: 12.68 gallons

Heater: On

Supplemental light: 4 hours (10pm-2am)

Harvest: No.

Summary of ScienceDome operations: GHO Tyler Hines’ experiment is continuing to be monitored. Forty LDEF seeds placed in the GreenHab are continuing to be monitored as well.

Dual split: On a timer from 10pm to 7am and under Mission Support supervision.

Summary of RAM operations: No operations were conducted in the RAM.

Summary of any observatory issues: The MDRS Robotic Observatory is inoperable until further notice due to a mechanical issue described in Sol 9’s Operations report.

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

Sol Summary Report – February 8th



Crew 274 Sol Summary Report 02-08-2023

Sol: 10

Summary Title: Save KURT

Author’s name: Sarah E. Guthrie (“Ceres”), Commander

Mission Status: Nominal

Sol Activity Summary: Two EVAs were completed with the remaining crew members for the rescue study utilizing the specialized rescue vest and medical sled for extracting incapacitated astronauts from the surface in locations not easily accessible by rovers, with special consideration for suit design. HSO Pender’s extra cache supplies at the Habitat (base supply) came in handy for the EVA crew as they were able to utilize the hydration packs for the strenuous EVAs today. Hab Specialist O’Hara continues his evaluation of habitats for this post-doctoral study and for Blue Origin. GHO Hines is monitoring the germination of the ldEF seeds which have surprisingly germinated after being exposed to the cosmic radiation of space for six years. Astronomers Pena and Loy are finally collecting night sky images of their HADS variable star after the NEW Mexico site was down for snowy weather. The crew looks forward to the last days of the mission.

Look Ahead Plan: The crew is closing up projects as this mission comes to a close and retrieving the cache on the last EVA of this mission (approval pending).

Anomalies in work: None

Weather: Sunny, 0C High, -3C Low; sunny

Crew Physical Status: Optimal

EVA: EVA #10 & 11 completed – successful

Reports to be filed: Sol Summary, Operations, EVA Request, EVA Report, GHO Report, and Astronomy Report.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site